OpenStudy (anonymous):

For a medal What is the value of the x variable in the solution to the following system of equations? 3x + y = 6 6x + 2y = 8 0 4 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system.

OpenStudy (anonymous):

:(

OpenStudy (ashleyisakitty):

There is no solution to this.

OpenStudy (anonymous):

can u help me with one more? please

OpenStudy (anonymous):

if the system is 1. a1x+b1y=c1 a2x+b2y=c2 \[if \frac{ a1 }{a2 }=\frac{ b1 }{ b2 }=\frac{ c1 }{ c2 },then system has infinite solutions\] \[if \frac{ a1 }{a2 }=\frac{ b1 }{ b2 }\neq \frac{ c1 }{c2 },then system has no solution.\]